### v1

# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

Abstract
Of concern is the Cauchy problem for the fractional semilinear heat equation with a forcing term
$\begin{cases}\partial_{t}u+(-\Delta)^{\frac{\theta}{2}}u=|u|^{\gamma-1}u+\mu,&\quad x\in\mathbb{R}^{N},\>t>0,\\ u(x,0)\equiv 0,&\quad x\in\mathbb{R}^{N},\end{cases}$ (P)
where $0<\theta\leqslant 2$, $\gamma>1,\>N\in\mathbb{Z}_{\geqslant 1}$ and $\mu$ is a tempered distribution on $\mathbb{R}^{N}$. In terms of Besov-type spaces, we give necessary conditions and sufficient conditions on $\mu$ for the local-in-time solvability of $\rm{(P)}$ in Lorentz spaces.
Address:
Graduate School of Mathematical Sciences, The University of Tokyo,
3-8-1 Komaba, Meguro-ku, Tokyo 153-8914, Japan.
E-mail: oka@ms.u-tokyo.ac.jp
2020 AMS Subject Classification. 35A01, 35K58, 35R11
Keywords. fractional heat equation, inhomogeneous Besov spaces, inhomogeneous term, solvability
1 Introduction and main results
We study the Cauchy problem for the fractional semilinear heat equation
$\begin{cases}\partial_{t}u+(-\Delta)^{\frac{\theta}{2}}u=|u|^{\gamma-1}u+\mu,&\quad\text{in }\mathbb{R}^{N},\>t>0,\\ u(x,0)\equiv 0,&\quad x\in\mathbb{R}^{N},\end{cases}$ (1.1)
where $\partial_{t}:=\partial/\partial t$, $\gamma>1$, $\theta\in\left]0,2\right]$, $N\in\mathbb{Z}_{\geqslant 1}$ and the forcing term $\mu=\mu(x)$ is a tempered distribution on $\mathbb{R}^{N}$. The symbol $(-\Delta)^{\frac{\theta}{2}}$ represents the fractional power of the Laplace operator $\Delta$ in $\mathbb{R}^{N}$, i.e.,
$(-\Delta)^{\frac{\theta}{2}}f(x)=\mathcal{F}^{-1}\left[|\xi|^{\theta}\mathcal{F}f(\xi)\right](x),\quad x\in\mathbb{R}^{N}.$
Here $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse on $\mathbb{R}^{N}$, respectively.
The aim of the present paper is twofold. First, by utilizing the framework of Besov-type spaces, we construct local-in-time solutions to problem (1.1) for a wide class of forcing terms, including distributions such as the derivative of Dirac delta, as well as Radon measures. Second, in terms of Besov-type spaces, we provide a necessary condition on the forcing term $\mu$ for the existence of local-in-time (possibly) sign-changing solutions belonging to an appropriate function space.
Problem (1.1) has been studied by many mathematicians (see e.g., [3, 9, 11, 12, 13, 14, 27, 28, 29] and references therein) from the perspective of solvability. In these works, attention has been paid to various characteristics of the forcing terms, such as their decay rate at spatial infinity and certain integral quantities. In particular, spatial singularities of forcing terms were considered in [9, 11] and some sufficient conditions and necessary conditions on the forcing terms are demonstrated from the viewpoints of local growth rates and function spaces.
We first focus on sufficient conditions for the existence of solutions to (1.1). Hisa, Ishige and Takahashi [9] aimed to identify the spatially strongest singular profile of $\mu$ for the existence of nonnegative solutions and succeeded in providing some solvable criteria for $\mu$. Their results concerning sufficient conditions can be summarized as follows. There exists a constant $A>0$ depending on $N$, $\theta$ and $\gamma$ such that the following claims hold.
  1. (1)
    Let $\gamma<\frac{N}{N-\theta}$. If a nonnegative Radon measure $\mu$ in $\mathbb{R}^{N}$ satisfies
    $\sup_{x\in\mathbb{R}^{N}}\mu(B(x,T^{1/\theta}))\leqslant AT^{\frac{N}{\theta}-\frac{\gamma}{\gamma-1}}$
    for some $T>0$, then problem (1.1) admits a solution in $\mathbb{R}^{N}\times\left[0,T\right[$.
  2. (2)
    Let $\gamma\geqslant\frac{N}{N-\theta}$. If a nonnegative measurable function $\mu:\mathbb{R}^{N}\to\mathbb{R}$ satisfies
    $0\leqslant\mu(x)\leqslant\begin{cases}A|x|^{-\frac{\theta\gamma}{\gamma-1}}+C_{0}&\text{if }\gamma>\frac{N}{N-\theta},\\ A|x|^{-N}\left[\log\left(e+\frac{1}{|x|}\right)\right]^{-\frac{N}{\theta}}+C_{0}&\text{if }\gamma=\frac{N}{N-\theta},\end{cases}$
    for some $C_{0}>0$, then problem (1.1) admits a local-in-time solution.
Theorem (1) is obtained by the contraction mapping theorem in uniformly local $L^{\gamma}$ spaces, while (2) is shown by constructing of some supersolutions.
Ishige, Kawakami and Takada [11] found sufficient conditions on $\mu$ for the case of $\gamma\geqslant\frac{N}{N-\theta}$ by combining the real interpolation method with the uniformly local Zygmund type spaces, which was recently introduced in [10] and applied to Fujita critical Cauchy problem ((1.4) below with $\gamma=1+\theta/N$). Their results were proved by fixed point theorem in uniformly local Lorentz spaces/Zygmund type spaces and were consistent with Theorem (2). For the definition of uniformly local Lorentz spaces, see Definition 1.1 (We do not work with Zygmund type spaces, see [10, 11] for detail). Their sufficient conditions are sharp, but unfortunately, their function spaces does not contain distributional forcing terms.
In this paper, we adopt uniformly local Lorentz-based Besov spaces (see Definition 1.3) to give sufficient conditions on the forcing term $\mu(x)$ for the existence of solutions (see Theorem 1.5). This may be regarded as an inhomogeneous problem version of the works by Kozono and Yamazaki [17] and by Zhanpeisov [30] which constructed local-in-time solutions to problem (1.4) for distributional initial data. In the case of $\gamma>\frac{N}{N-\theta}$, our result covers the previous studies [9, 11] and our result gives a slightly looser condition on $\mu$ than theirs (see Remark 1.6 and Remark 1.7). Our existence theorem also works well in the case of $\gamma<\frac{N}{N-\theta}$; the derivative of Dirac delta is dealt under suitable conditions on $\theta$ and $\gamma$, as well as the sufficient condition in [9] is almost covered (see Remark 1.8).
Secondly, we tackle necessary conditions for the existence of solutions. Hisa, Ishige and Takahashi [9] (building on [8, 13, 14]) found a necessary condition on Radon measure $\mu$ for the existence of nonnegative local-in-time solutions to (1.1). If $u$ is a nonnegative solution to problem (1.1) on $\left]0,T\right[$ for some $T>0$, then the inhomogeneous term $\mu$ is a Radon measure and satisfies the following growth conditions: There exist constants $A_{0}$, $A_{1}>0$ such that the estimate
$\sup_{x\in\mathbb{R}^{N}}\mu(B(x,\sigma))\leqslant A_{0}\sigma^{N-\frac{\theta\gamma}{\gamma-1}}\quad\text{for }\sigma\in\left]0,T^{1/\theta}\right]$
holds. In addition, when $\gamma=\frac{N}{N-\theta}$, it holds that
$\sup_{x\in\mathbb{R}^{N}}\mu(B(x,\sigma))\leqslant A_{1}\left[\log\left(e+\frac{T^{1/\theta}}{\sigma}\right)\right]^{1-\frac{N}{\theta}}\quad\text{for }\sigma\in\left]0,T^{1/\theta}\right].$
These conditions are sharp, considering their sufficient conditions (Claims (1) and (2)). We remark that their argument strongly depends on the nonnegative assumption on $u$ and thus we cannot apply their theory to sign-changing solutions.
Recently, in the study of the Cauchy problem of the Navier–Stokes equation in $\mathbb{R}^{N}$, Kozono, Okada and Shimizu [15, 16] proved that if there exists a Serrin class solution, then the initial data necessarily belongs to a specific homogeneous/inhomogeneous Besov spaces $B^{-1+N/p}_{p,q}$ (They also proved the existence theorem of Serrin class solutions for initial data which belong to $B^{-1+N/p}_{p,q}$). They used the characterization of homogeneous/inhomogeneous Besov spaces via heat semigroup (see e.g., [2, 5, 25]). It is noteworthy that positivity of $u$ is not assumed in their results.
Inspired by their works, we obtain necessary conditions on the forcing term $\mu$ for the existence of solutions to (1.1) in Lorentz spaces (see Theorem 1.9). Although our approach requires solutions to belong to some function spaces, which seems to be a strong assumption compared to [9], we do not need to impose the positivity assumption on $u$. This feature serves as a comparative advantage over [9]. Since the manner in which distributions $\mu$ appear in (1.1) is substantially different from that in (1.4) (cf. (1.2) and (1.5)), a new estimate for the Besov-type norm of $\mu$ needs to be established (see Proposition 2.16).
Throughout the following, the letter $C$ denotes a generic positive constant, which may change from line to line.
1.1 Function spaces and definition of solutions
We introduce some function spaces. Henceforth, the symbol $\left\|f\mid X\right\|$ denotes the norm of $f$ in any normed space $X$.
First we define the spaces for our solutions to problem (1.1). For any mesurable function $f$, we define a map $\alpha_{f}:\mathbb{R}_{>0}\to\left[0,\infty\right]$ as
$\alpha_{f}(\sigma):=\left|\left\{x\in\mathbb{R}^{N}\mid|f(x)|>\sigma\right\}\right|$
and a function $f^{\ast}:\mathbb{R}_{>0}\to\left[0,\infty\right]$ as
$f^{\ast}(\lambda)$ $:=\sup\left\{\sigma\in\mathbb{R}_{>0}\mathrel{}\middle|\mathrel{}\alpha_{f}(\sigma)>\lambda\right\}$
, where $|E|$ denotes the $N$-dimensional Lebesgue measure for any measurable set $E$. Here we adopt the convention $\sup\varnothing=0$.
Definition 1.1.
(Lorentz spaces)
  1. 1.
    Let $p\in\left]0,\infty\right[$ and $q\in\left]0,\infty\right]$. The Lorentz space $L^{p,q}(\mathbb{R}^{N})$ is defined as the set of measurable functions $f:\mathbb{R}^{N}\to\mathbb{R}$ such that $\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}\right\|<\infty$,
    $\text{where }\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}\right\|:=\begin{dcases}\left(\int_{0}^{\infty}\left(\lambda^{\frac{1}{p}}f^{\ast}(\lambda)\right)^{q}\,\frac{d\lambda}{\lambda}\right)^{1/q}&\text{for }q\in\left]0,\infty\right[,\\ \sup_{\lambda>0}\,\lambda^{\frac{1}{p}}f^{\ast}(\lambda)=\sup_{\lambda>0}\lambda\alpha_{f}(\lambda)^{1/p}&\text{for }q=\infty.\end{dcases}$
  2. 2.
    Let $p\in\left]0,\infty\right[$ and $q\in\left]0,\infty\right]$. The uniformly local Lorentz space $L^{p,q}_{\text{ul}}(\mathbb{R}^{N})$ is defined as the set of measurable functions $f:\mathbb{R}^{N}\to\mathbb{R}$ such that
    $\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}_{\mathrm{ul}}\right\|:=\sup_{z\in\mathbb{R}^{N}}\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|<\infty,$
    where $\chi_{E}$ is the characteristic function of $E(\subset\mathbb{R})$ and $B(x,r)$ is the open ball in $\mathbb{R}^{N}$ with center $x\in\mathbb{R}^{N}$ and radius $r>0$.
Remark 1.2.
  1. 1.
    Lorentz spaces $L^{p,q}$ are just quasi-normed spaces, but for $p\in\left]1,\infty\right[$ and $q\in\left[1,\infty\right]$, they can be equipped with an equivalent norm (see Remark 2.7). In this paper, we only use these spaces in the case of $p>\gamma>1$, $q\geqslant 1$ and thus we regard the spaces $\left(L^{p,q},\left\|\cdot\mathrel{}\middle|\mathrel{}L^{p,q}\right\|\right)$ as Banach spaces by abuse of notation.
  2. 2.
    Choosing $p$ and $q$ so that the Lorentz space $L^{p,q}$ is a Banach space, we can prove that the uniformly local Lorentz space $L^{p,q}_{\mathrm{ul}}$ is also a Banach space.
We introduce the inhomogeneous Littlewood–Paley dyadic decomposition in order to define the inhomogeneous Besov-type spaces. One can consult [24] for a general reference. Let $\zeta(t)\in C_{c}^{\infty}\left(\left[0,\infty\right[;\mathbb{R}\right)\>$ satisfy  $0\leqslant\zeta(t)\leqslant 1,\>\zeta(t)\equiv 1$ for $t\leqslant 3/2$ and $\text{supp}\>\zeta\subset\left[0,5/3\right[$. We set
$\varphi_{(0)}(\xi):=\zeta(|\xi|),\quad\varphi_{0}(\xi):=\zeta(|\xi|)-\zeta(2|\xi|),\quad\varphi_{j}(\xi):=\varphi_{0}(2^{-j}\xi)\quad\text{for}\,\,j\in\mathbb{Z}.$
Then, we have $\varphi_{(0)}(\xi),\varphi_{j}(\xi)\in C_{c}^{\infty}(\mathbb{R}^{N})$ for all $j\in\mathbb{Z}$ and
$\varphi_{(0)}(\xi)+\sum_{j=1}^{\infty}\varphi_{j}(\xi)=\lim_{n\rightarrow\infty}\zeta(2^{-n}|\xi|)=1\quad\text{for all }\xi\in\mathbb{R}^{N}.$
[…]
$\Delta_{j}f:=\begin{cases}\mathcal{F}^{-1}\left[\varphi_{(0)}(\xi)\mathcal{F}f\right]&\quad\text{for }j=0,\\ \mathcal{F}^{-1}\left[\varphi_{j}(\xi)\mathcal{F}f\right]&\quad\text{for }j\geqslant 1.\end{cases}$
Now we define the spaces of our forcing terms.
Definition 1.3.
(inhomogeneous Besov–Lorentz spaces) Let $1<p<\infty$, $q,r\in\left[1,\infty\right]$ and $s\in\mathbb{R}$. The inhomogeneous Besov–Lorentz space $B^{s}_{(p,q),r}(\mathbb{R}^{N})$ is defined as the set of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ such that
$\Delta_{j}f$ $\in L^{p,q}_{\mathrm{ul}}(\mathbb{R}^{N})\quad\text{for all }j\in\mathbb{N}\quad\text{and }$
$\left\|f\mid B^{s}_{(p,q),r}\right\|$ $\coloneqq\left\|\left\{2^{sj}\left\|\Delta_{j}f\mid L^{p,q}_{\mathrm{ul}}\right\|\right\}_{j\in\mathbb{N}}\mathrel{}\middle|\mathrel{}l^{r}(\mathbb{N})\right\|<\infty.$
This kind of generalization of Besov spaces was introduced by [17] and has been effectively used in the study of semilinear parabolic equations (see e.g., [22, 30]). Their general properties can be found in [19]. The reason we do not employ Morrey spaces but instead employ Lorentz spaces as a base of Besov spaces lies in the Serrin supercritical case (see Lemma 3.3). We mainly use spaces $B^{s}_{(p,q),r}$ for the case of $q$, $r\in\left\{1,\infty\right\}$. Next we put $S(t)f:=\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\mathcal{F}f\right]$ for every $f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and $t>0$ to formulate our definition of solutions.
Definition 1.4.
Let $1<a<\infty$, $s\in\mathbb{R}$, $q,r\in\left[1,\infty\right]$ and $T>0$. For $\mu\in B^{s}_{(a,q),r}(\mathbb{R}^{N})$, we say that a measurable function $u:\mathbb{R}^{N}\times\left]0,T\right[\to\mathbb{R}$ is a solution to (1.1) on $\left]0,T\right[$ if $u$ satisfies
$u\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$
for some $1<p<\infty$ and the following integral equation
$u(x,t)=\int_{0}^{t}S\left(t-\tau\right)\left[\left|u\right|^{\gamma-1}u(\cdot,\tau)\right](x)\,d\tau+\int_{0}^{t}S\left(t-\tau\right)\mu(x)\,d\tau$ (1.2)
for a.e.-$(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$.
There is a possibility that the second term of the right hand side of (1.2) is not a function if $\mu$ is a general tempered distribution. In our existence theorem (Theorem 1.5), such cases are eliminated and it certainly exists as an element of $L^{\infty}\left(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right)$. We do not consider about initial condition in Definition 1.4 but we make a comment on it (see Remark 3.5).
1.2 Main results
First we state sufficient conditions for local-in-time existence.
Theorem 1.5.
Let $\gamma>1$, $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$.
  1. (1)
    There exist positive constants $\delta$, $M$ and $T$ such that for every $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{\left(p,\infty\right),1}\right\|<\delta$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ with a bound
    $\sup_{t\in\left]0,T\right[}\left\|u(x,t)\mid L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right\|\leqslant M.$ (1.3)
  2. (2)
    There exist $\delta>0$ and $M>0$ such that for every $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying
    $\mu\in B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}(\mathbb{R}^{N})\quad\text{and}\quad\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right\|<\delta$
    for some $\varepsilon\in\left]0,\theta\right[$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.3).
  3. (3)
    Assume $\theta<s<0$. There exists $M>0$ such that for every $\mu\in B^{s}_{(p,\infty),\infty}(\mathbb{R}^{N})$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.3). ∎
Each of these three sufficient conditions on $\mu$ has its own advantages. Note that $p=\frac{N(\gamma-1)}{\theta}$ is feasible only when the Serrin supercritical condition $\gamma>\frac{N}{N-\theta}$ holds, because we imposed $p>\gamma$.
Remark 1.6.
From Proposition 2.9 and Proposition 2.10, we have
$L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}\subset B^{0}_{\left(\frac{N(\gamma-1)}{\gamma\theta},\infty\right),\infty}\subset B^{\varepsilon-\theta}_{\left(\frac{N(\gamma-1)}{\theta+(\gamma-1)\varepsilon},\infty\right),\infty}$
for some $\varepsilon>0$. Thus in the Serrin supercritical case $\gamma>\frac{N}{N-\theta}$, by applying Theorem 1.5(2) with $p=\frac{N(\gamma-1)}{\theta}$, we find out that problem (1.1) possesses a solution for $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}\right\|<\delta$. For example, we can take homogeneous functions $c|x|^{-\frac{\gamma\theta}{\gamma-1}}$ or their infinite sum
$c\sum_{j\in\mathbb{N}}|x-x_{j}|^{-\frac{\gamma\theta}{\gamma-1}}\chi_{B(x_{j},1)}(x),\quad|x_{j}-x_{k}|>1\text{ for }j\neq k$
as forcing terms for sufficiently small $c>0$. This is compatible with the result of [11]. Moreover, the embeddings above are non-trivial, our sufficient condition is slightly broader than their results.
Remark 1.7.
We know that all $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta}}(\mathbb{R}^{N})$ are acceptable as forcing term. Recall that problem (1.1) has the following structure: If a funciton $u(x,t)$ is a solution to problem (1.1) with a forcing term $\mu(x)$, then a function $u_{\lambda}(x,t):=\lambda^{\frac{\theta}{\gamma-1}}u(\lambda x,\lambda^{\theta}t)$ is a solution to problem (1.1) with $\mu_{\lambda}(x):=\lambda^{\frac{\gamma\theta}{\gamma-1}}\mu(\lambda x)$. For any $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta}}(\mathbb{R}^{N})$, we see that
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu_{\lambda}\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|$ $\leqslant\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{0}_{\left(\frac{N(\gamma-1)}{\gamma\theta},\infty\right),\infty}\right\|$
$\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\text{ul}}\right\|$ $\leqslant C\sup_{z\in\mathbb{R}^{N}}\left(\int_{|x-z|\leqslant 1}\lambda^{N}|\mu(\lambda x)|^{\frac{N(\gamma-1)}{\gamma\theta}}\,dx\right)^{\frac{\gamma\theta}{N(\gamma-1)}}$
$=C\sup_{z\in\mathbb{R}^{N}}\left(\int_{|x-z|\leqslant\lambda}|\mu(x)|^{\frac{N(\gamma-1)}{\gamma\theta}}\,dx\right)^{\frac{\gamma\theta}{N(\gamma-1)}}$
holds for all $\lambda>0$. By absolute continuity of $\left|\mu\right|^{\frac{N(\gamma-1)}{\gamma\theta}}$, the last term tends to zero if we send $\lambda\to 0$. Thus we can apply Remark 1.6 to $\mu_{\lambda}$, $0<\lambda\ll 1$.
Remark 1.8.
In the Serrin subcritical case $\gamma<\frac{N}{N-\theta}$, we can pick $p_{0}$ as $\gamma<p_{0}<\frac{N}{N-\theta}$. Then by Sobolev embedding (Proposition 2.10), we have
$L^{1}\subset B^{0}_{1,\infty}\subset B^{N/p_{0}-N}_{p_{0},\infty}\subset B^{N/p_{0}-N}_{(p_{0},\infty),\infty}$
and the space appearing in the last satisfies the assumptions of Theorem 1.5 (3). For instance, we know that the Dirac delta measure $\delta(x)$ belongs to $B^{0}_{1,\infty}\subset B^{N/p_{0}-N}_{p_{0},\infty}\subset B^{N/p_{0}-N}_{(p_{0},\infty),\infty}$. Furthermore, if $\theta\in\left]1,2\right]$ and $\gamma>1$ satisfies $\gamma<\frac{N}{N+1-\theta}$, we can take $\partial_{x_{i}}\delta(x)$ as a forcing term of problem (1.1) since we can take $p_{1}\in\left]\gamma,\frac{N}{N+1-\theta}\right[$ and we have
$\partial_{x_{i}}\delta(x)\in B^{N/p_{1}-N-1}_{p_{1},\infty}\subset B^{N/p_{1}-N-1}_{(p_{1},\infty),\infty}\quad\text{for}\quad i\in\left\{1,...,N\right\}.$
This example reflects a benefit of employing Besov-type spaces. We also note that employing Besov–Morrey spaces (see Definition 1.2 in [30]) can broaden the class of acceptable forcing terms in this case ($\gamma<N/(N-\theta)$).
In this paper, we do not treat in detail the Serrin critical case $\gamma=\frac{N}{N-\theta}$. This case requires some special function spaces (see Theorem 1.3 in [11] for details).
The next theorem states a necessary condition on forcing terms $\mu$ for the existence of a solution $u=u(x,t)$ belonging to $L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$.
Theorem 1.9.
Let $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ satisfies the integral equation (1.2) for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. Then, the tempered distribution $\mu$ appering in (1.2) must be an element of $B^{-\theta}_{(p,\infty),\infty}(\mathbb{R}^{N})$. ∎
Remark 1.10.
We claim nothing about existence of solutions to problem (1.1) for $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying
[…]
By similarly carrying the arguments of the proof of Theorem 1.9 with the aid of Theorem 2 in [1] and of the Hardy–Littlewood–Sobolev inequality (see also [16]), we can obtain a necessary condition on the solvability of the following Cauchy problem
$\begin{cases}\partial_{t}u+(-\Delta)^{\theta/2}u=|u|^{\gamma-1}u,&\quad x\in\mathbb{R}^{N},\>t>0,\\ u(x,0)=\mu(x)\in\mathcal{S}^{\prime}(\mathbb{R}^{N}),&\quad x\in\mathbb{R}^{N},\end{cases}$ (1.4)
where $\gamma>1$, $\theta\in\left]0,2\right]$.
Corollary 1.11.
Let $p$ and $q$ satisfy the conditions
$\frac{N(\gamma-1)}{\theta}<p<\infty,\quad\gamma\leqslant p,\quad\gamma<q<\infty\quad\text{and}\quad\frac{\theta}{\gamma-1}-\frac{\theta}{q}-\frac{N}{p}=0.$
Assume that a function $u(x,t)\in L^{q}(0,T;L^{p}(\mathbb{R}^{N}))$ satisfies the integral equation
$u(x,t)=S({t})\mu(x)+\int_{0}^{t}S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right](x)\,d\tau$ (1.5)
for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$ for some $T>0$. Then, the tempered distribution $\mu(x)$ appering in (1.5) must be an element of inhomogeneous Besov space $B^{-\theta/q}_{p,q}(\mathbb{R}^{N})=B^{N/p-\theta/(\gamma-1)}_{p,\theta p(\gamma-1)/(\theta p-N\gamma+N)}(\mathbb{R}^{N})$. ∎
The rest of this paper proceeds as follows. We divide Section 2 into five subsections. The first three are devoted to collecting basic facts, including real interpolation for Lorentz and Besov spaces. We recall decay estimates of operators $S(t)$ in the next subsection. At the end of Section 2, we make an estimate of the Besov–Lorentz norm of the forcing term. When $\theta<2$, it needs additional care and we postpone this task to Appendix A. In Section 3, we prove Theorem 1.5 by applying the contraction mapping theorem and Theorem 1.9 by applying Proposition 2.16.
2 Preliminaries
In this section, we gather some properties of Lorentz spaces $L^{p,q}$, $L^{p,q}_{\mathrm{ul}}$, of Besov–Lorentz spaces $B^{s}_{(p,q),r}$ and of operators $S(t)$ on these spaces. We also show that Besov–Lorentz norms $\left\|f\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,q),\infty}\right\|$ can be estimated by a space-time norm of the Duhamel integral of $f$ (Proposition 2.16).
2.1 Basic properties of Lorentz spaces
We begin with elemental properties of Lorentz spaces.
Lemma 2.1.
Fix $\mathcal{L}\in\left\{L,L_{\mathrm{ul}}\right\}$ and $p\in\left]0,\infty\right[$. The following claims hold.
  1. (1)
    Assume that $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p^{-1}=p_{0}^{-1}+p_{1}^{-1}$. Then, there exists $C>0$ depending on $p_{0}$, $p_{1}$ such that the inequality
    $\left\|fg\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|\leqslant C\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{p_{0},\infty}\right\|\left\|g\mathrel{}\middle|\mathrel{}\mathcal{L}^{p_{1},\infty}\right\|$
    holds for all $f\in\mathcal{L}^{p_{0},\infty}$ and $g\in\mathcal{L}^{p_{1},\infty}$.
  2. (2)
    Fix $r\in\left]0,\infty\right[$. Then, the relation
    $\left\||f|^{r}\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|=\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{pr,\infty}\right\|^{r}$
    holds for all $f\in\mathcal{L}^{pr,\infty}$.
  3. (3)
    Assume $|g(x)|\leqslant|f(x)|$ for a.e.-$x$. Then, we have
    $\left\|g\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|.$
  4. (4)
    We have the following scaling property. For all $\lambda>0$ and $f\in L^{p,q}$, we have
    $\left\|f_{\lambda}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|=\lambda^{\frac{-N}{p}}\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}\right\|$
    , where $f_{\lambda}(x):=f(\lambda x)$ and $q\in\left]0,\infty\right[$.
Proof.
(1) for $\mathcal{L}=L$: See Exercise 1.1.15 in [7]. Others are easily confirmed. ∎
We recall the Young inequality for Lorentz spaces.
Proposition 2.2.
Let $p\in\left]1,\infty\right[$, $q\in\left[1,\infty\right]$. There exists a positive number $C$ depending on $p$ such that the inequality
$\left\|g\ast f\mid L^{p,q}(\mathbb{R}^{N})\right\|\leqslant C\left\|g\mid L^{1}(\mathbb{R}^{N})\right\|\left\|f\mid L^{p,q}(\mathbb{R}^{N})\right\|$
holds for all $g\in L^{1}(\mathbb{R}^{N})$ and $f\in L^{p,\infty}(\mathbb{R}^{N})$.
Proof.
See Theorem 2.6. in [23] (For $q=\infty$, it is enough to see Theorem 1.2.13 in [7] (pp. 23)). ∎
We prepare the Young inequality for uniformly local Lorentz spaces.
Proposition 2.3.
Let $p\in\left]1,\infty\right[$. There exists a positive constant $C$ depending on $p$ such that the inequality
$\left\|g\ast f\mid L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})\right\|\leqslant C\left\|g\mid L^{1}(\mathbb{R}^{N})\right\|\left\|f\mid L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})\right\|$
holds for all $g\in L^{1}(\mathbb{R}^{N})$ and $f\in L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})$.
It is worth mentioning that Propostion 2.3 does not directly follow from Proposition 2.2. We shall modify the proof of Proposition 2.2 (see pp. 24–25 in [7]) and provide the full proof here for completeness.
Proof of Proposition 2.3.
Fix $z\in\mathbb{R}^{N}$ and $\lambda>0$. Take any $f\in L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})$ and $g\in L^{1}(\mathbb{R}^{N})$. Set $M:=\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}$. We decompose $f$ as
$f(x)=f(x)\chi_{\left\{|f(x)|\leqslant M\right\}}(x)+f(x)\chi_{\left\{|f(x)|>M\right\}}(x)=:f_{0}(x)+f_{1}(x).$
Because $\left|f_{0}\ast g(x)\right|\leqslant\lambda/2$ holds for all $x\in\mathbb{R}^{N}$, we have
$\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)\leqslant\alpha_{(f_{0}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)+\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)=\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right).$
By Chebyshev’s inequality, we get
$\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)$ $\leqslant\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)\leqslant\frac{2}{\lambda}\left\|f_{1}\ast g\mathrel{}\middle|\mathrel{}L^{1}(B(z,1))\right\|$
$\leqslant\frac{2}{\lambda}\int_{B(z,1)}\int_{\mathbb{R}^{N}}\left|f_{1}(x-y)\right||g(y)|\,dydx$
$=\frac{2}{\lambda}\int_{\mathbb{R}^{N}}|g(y)|\int_{B(z+y,1)}\left|f_{1}(x)\right|\,dxdy.$ (2.1)
Here, for fixed $y\in\mathbb{R}^{N}$, we carry out the following calculation.
$\int_{B(z+y,1)}\left|f_{1}(x)\right|\,dx$ $=\int_{B(z+y,1)}\int_{0}^{\infty}\chi_{\left\{0<\lambda<|f_{1}(x)|\right\}}(\lambda)\,d\lambda dx=\int_{0}^{\infty}\alpha_{f_{1}\chi_{B(z+y,1)}}(\lambda)\,d\lambda$
$=\int_{0}^{M}\alpha_{f\chi_{B(z+y,1)}}(M)\,d\lambda+\int_{M}^{\infty}\alpha_{f\chi_{B(z+y,1)}}(\lambda)\lambda^{p}\lambda^{-p}\,d\lambda$
$\leqslant M^{1-p}\left\|f\chi_{B(z+y,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{p}+\int_{M}^{\infty}\frac{d\lambda}{\lambda^{p}}\left\|f\chi_{B(z+y,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{p}$
$\leqslant\frac{p}{p-1}M^{1-p}\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{p}.$ (2.2)
We plug (2.1) into (2.1) and recall $M=\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}$ to see that
$\lambda\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)^{1/p}$ $\leqslant\lambda\left[\frac{2}{\lambda}\int_{\mathbb{R}^{N}}|g(y)|\frac{p}{p-1}\left(\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}\right)^{1-p}\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{p}\,dy\right]^{1/p}$
$\leqslant 2\left(\frac{p}{p-1}\right)^{1/p}\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
holds for all $\lambda>0$ and $z\in\mathbb{R}^{N}$. Taking the sup in $\lambda$ and $z$ yields the claim. ∎
2.2 Real interpolation of Lorentz spaces
Next we introduce some notations for real interpolation. We can pick the following from [4].
Definition 2.4.
Let $X_{0}$ and $X_{1}$ be quasi-normed spaces and suppose that there exists a Hausdorff topological vector space $\mathfrak{X}$ such that $X_{0}$ and $X_{1}$ are subspaces of $\mathfrak{X}$. For $f\in X_{0}+X_{1}$ and $\lambda>0$, we put
$K(f,\lambda;X_{0},X_{1}):=\inf_{\begin{subarray}{c}f=f_{0}+f_{1}\\ f_{0}\in X_{0},\,f_{1}\in X_{1}\end{subarray}}\big\{\left\|f_{0}\mathrel{}\middle|\mathrel{}X_{0}\right\|+\lambda\left\|f_{1}\mathrel{}\middle|\mathrel{}X_{1}\right\|\big\}.$
Then for $\kappa\in\left]0,1\right[$ and $q\in\left]0,\infty\right]$, we define the space $\left(X_{0},X_{1}\right)_{\kappa,q}$ as
$\left(X_{0},X_{1}\right)_{\kappa,q}:=\left\{f\in X_{0}+X_{1}\mathrel{}\middle|\mathrel{}\left\|f\mathrel{}\middle|\mathrel{}\left(X_{0},X_{1}\right)_{\kappa,q}\right\|<\infty\right\}$
, where
$\left\|f\mathrel{}\middle|\mathrel{}\left(X_{0},X_{1}\right)_{\kappa,q}\right\|:=\begin{dcases}\left(\int_{0}^{\infty}\left(\lambda^{-\kappa}K(f,\lambda;X_{0},X_{1})\right)^{q}\,\frac{d\lambda}{\lambda}\right)^{1/q}&\text{for }q\in\left]0,\infty\right[,\\ \sup_{\lambda>0}\,\lambda^{-\kappa}K(f,\lambda;X_{0},X_{1})&\text{for }q=\infty.\end{dcases}$
Proposition 2.5.
Take $X_{0}$ and $X_{1}$ as in Definition 2.4. Let $\kappa\in\left]0,1\right[$, $q\in\left]0,\infty\right]$.
  1. (1)
    If $X_{0}$ and $X_{1}$ are normed spaces and $q\geqslant 1$, then $(X_{0},X_{1})_{\kappa,q}$ is a normed space.
  2. (2)
    If $X_{0}$ and $X_{1}$ are complete, then $(X_{0},X_{1})_{\kappa,q}$ is complete.
Proof.
(1)   It can be easily checked.  (2)   See Theorem 3.4.2 in [4]. ∎
Theorem 2.6.
Let $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p_{0}\neq p_{1}$ and take $q\in\left]0,\infty\right]$ and $\kappa\in\left]0,1\right[$. We set $p\in\left]0,\infty\right[$ by the formula $p^{-1}=(1-\kappa)p_{0}^{-1}+\kappa p_{1}^{-1}$. Then, we have $\left(L^{p_{0},q_{0}},L^{p_{1},q_{1}}\right)_{\kappa,q}=L^{p,q}$ for all $q_{0},q_{1}\in\left]0,\infty\right]$.
Proof.
See Theorem 5.3.1 in [4]. ∎
Remark 2.7.
If $1<p<\infty$ and $1\leqslant q\leqslant\infty$, we may write the space $L^{p,q}$ as
$L^{p,q}=\left(L^{p_{0},p_{0}},L^{p_{1},p_{1}}\right)_{\kappa,q}=\left(L^{p_{0}},L^{p_{1}}\right)_{\kappa,q}$
with $p_{0},p_{1}\in\left]1,\infty\right[$ by means of Theorem 2.6. Thus, Proposition 2.5 (1) and (2) reveal the space $L^{p,q}$ as a Banach space.
It is useful for Section 3 to prepare a $L^{p,q}_{\mathrm{ul}}$ version of Theorem 2.6.
Corollary 2.8.
Let $p_{0}$, $p_{1}$, $q$, $\kappa$ and $p$ be as in Theorem 2.6. Then, the embedding
$\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}\subset L^{p,q}_{\mathrm{ul}}$
is valid for all $q_{0}$, $q_{1}\in\left]0,\infty\right]$.
Proof.
Fix $z\in\mathbb{R}^{N}$ and $\lambda>0$. Take $f\in\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}$. Define $A$ and $B(\subset\mathbb{R})$ as
$A(z,\lambda,f)$ $:=\left\{\left\|f_{0}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}\right\|+\lambda\left\|f_{1}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f\chi_{B(z,1)}=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}},\,\,f_{1}\in L^{p_{1},q_{1}}\end{subarray}\right\},$
$B(z,\lambda,f)$ $:=\left\{\left\|f_{0}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}\right\|+\lambda\left\|f_{1}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}}_{\mathrm{ul}},\qquad f_{1}\in L^{p_{1},q_{1}}_{\mathrm{ul}}\end{subarray}\right\}.$
We can easily check $A(z,\lambda,f)\supset B(z,\lambda,f)$ and this inclusion implies that
$K(f\chi_{B(z,1)},\lambda;L^{p_{0},q_{0}},L^{p_{1},q_{1}})=\inf A(z,\lambda,f)\leqslant\inf B(z,\lambda,f)$
$\leqslant\inf\left\{\left\|f_{0}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}_{\mathrm{ul}}\right\|+\lambda\left\|f_{1}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}_{\mathrm{ul}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}}_{\mathrm{ul}},\qquad f_{1}\in L^{p_{1},q_{1}}_{\mathrm{ul}}\end{subarray}\right\}=K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})$
for all $z\in\mathbb{R}^{N}$ and $\lambda>0$. In the light of Theorem 2.6, we have
$\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|$ $\leqslant C\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}\left(L^{p_{0},q_{0}},L^{p_{1},q_{1}}\right)_{\kappa,q}\right\|$
$\leqslant C\begin{dcases}\left(\int_{0}^{\infty}\left(\lambda^{-\kappa}K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})\right)^{q}\,\frac{d\lambda}{\lambda}\right)^{1/q}&\text{for }q\in\left]0,\infty\right[,\\ \sup_{\lambda>0}\,\lambda^{-\kappa}K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})&\text{for }q=\infty\end{dcases}$
$=C\left\|f\mathrel{}\middle|\mathrel{}\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}\right\|.$
Taking the supremum over $z\in\mathbb{R}^{N}$ leads the desired embedding. ∎
2.3 Properties of Besov–Lorentz spaces
We prove some properties of Besov–Lorentz spaces, with reference to [17].
Proposition 2.9.
We have the following continuous embeddings
$B^{0}_{(p,\infty),1}\subset L^{p,\infty}_{\mathrm{ul}}\subset B^{0}_{(p,\infty),\infty}$
for $p\in\left]1,\infty\right[$.
Proof.
By using Proposition 2.3, we have
$\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\begin{cases}C\left\|\mathcal{F}^{-1}\varphi_{j}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|=C\left\|\mathcal{F}^{-1}\varphi_{0}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|&\text{for }j\geqslant 1\\ C\left\|\mathcal{F}^{-1}\varphi_{(0)}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|&\text{for }j=0\end{cases}$
for all $f\in L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})$. Taking the supremum over $j\in\mathbb{N}$ gives the second inclusion. Next we take $f\in B^{0}_{(p,\infty),1}$. We can easily check that $f=\sum_{j\in\mathbb{N}}\Delta_{j}f$ in $\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and can conduct the following calculation
$\left\|\sum_{j=0}^{n}\Delta_{j}f-\sum_{j=0}^{m}\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\sum_{j=m+1}^{n}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\left(\leqslant\left\|f\mid B^{0}_{(p,\infty),1}\right\|<\infty\right)$
$\rightarrow 0\quad(\text{as }m,n\to\infty).$
Thus $\sum_{j\in\mathbb{N}}\Delta_{j}f=f$ belongs to $L^{p,\infty}_{\mathrm{ul}}$. The continuity is easily seen. ∎
Note that we didn’t use properties of Lorentz spaces apart from their completeness and Proposition 2.3 (see also Proposition 2.11 in [17]). We next verify the embedding theorem of Sobolev-type.
Proposition 2.10.
(Sobolev-type embedding) Let $1\leqslant r\leqslant\infty$, $s\in\mathbb{R}$, $1<p<\infty$ and $1\leqslant q\leqslant\infty$. We have the following continuous embeddings:
  1. (1)
    $B^{s}_{(p,q),r}\subset B^{s-N/p}_{\infty,r}$.
  2. (2)
    $B^{s}_{(p,q),r}\subset B^{s-N(1-\theta)/p}_{\left(p/\theta,q/\theta\right),r}$ for $\theta\in\left]0,1\right[$.
Here, $B^{s-N/p}_{P,r}$ stands for inhomogeneous Besov spaces defined in the same way as inhomogeneous Besov–Lorentz spaces by replacing Lorentz spaces with standard $L^{P}$ space.
To obtain this, we modify the proof of Theorem 2.5 in [17]. We mention that there is another way to prove this claim (see Remark 2.11 below).
Proof of Proposition 2.10.
First we show the following inequality
$\sup_{x\in\mathbb{R}^{N}}\left|\Delta_{j}f(x)\right|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mid L^{p,q}_{\mathrm{ul}}\right\|$ (2.3)
is valid for all $f\in L^{p,q}_{\mathrm{ul}}$ and for all $j\in\mathbb{N}$ (the latter is a standard fact).
We write $\Phi_{j}(\xi):=\varphi_{j-1}(\xi)+\varphi_{j}(\xi)+\varphi_{j+1}(\xi)$ for $j\in\mathbb{Z}$ and $\Phi_{(0)}(\xi):=\varphi_{(0)}(\xi)+\varphi_{1}(\xi)$. Then identities $\varphi_{j}(\xi)\equiv\Phi_{j}(\xi)\varphi_{j}(\xi)$, $\varphi_{(0)}(\xi)\equiv\Phi_{(0)}(\xi)\varphi_{(0)}(\xi)$ and $\Psi_{j}(x):=\mathcal{F}^{-1}\Phi_{j}(x)=2^{Nj}\mathcal{F}^{-1}\Phi_{0}(2^{j}x)$ hold. Fix $x\in\mathbb{R}^{N}$. We decompose $|\Delta_{j}f(x)|$ as
$\left|\Delta_{j}f(x)\right|$ $=\left|\mathcal{F}^{-1}\left[\Phi_{j}\varphi_{j}\mathcal{F}f\right](x)\right|\leqslant\int_{\mathbb{R}^{N}}\left|\Psi_{j}(y)\right|\left|\Delta_{j}f(x-y)\right|\,dy$
$\leqslant 2^{Nj}\left(\int_{B(0,1)}+\sum_{k=0}^{\infty}\int_{B(0,2^{k+1})\setminus B(0,2^{k})}\right)\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy.$ (2.4)
For the first term in (2.3), we deduce that
$2^{Nj}\int_{B(0,1)}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy\leqslant 2^{Nj}\int_{\mathbb{R}^{N}}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(y)\chi_{B(x,1)}(y)\right|\,dy$
$\leqslant$  $\,2^{Nj}\left\|\Psi_{0}(2^{j}y)\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|\leqslant 2^{Nj/p}\left\|\Psi_{0}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$.
Note that there exists a positive integer $\Omega=\Omega_{N}$ such that
$B(0,2)\subset\bigcup_{m=1}^{\Omega}B(z_{m},1)\quad\text{for some}\quad\left\{z_{m}\right\}_{m=1}^{\Omega}$
and therefore we can pick $\left\{z_{m}\right\}_{m=1}^{\Omega^{k}}$ satisfying $B(0,2^{k})\subset\bigcup_{m=1}^{\Omega^{k}}B(z_{m},1)$ for each $k\in\mathbb{N}$. Since $\Psi_{0}=\mathcal{F}^{-1}\Phi_{0}$ is a rapidly decreasing function, it holds that
$2^{Nj}\int_{B(0,2^{k+1})\setminus B(0,2^{k})}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy$
$\leqslant 2^{Nj}\sum_{m=1}^{\Omega^{k+1}}\int_{\mathbb{R}^{N}}\left|\Psi_{0}(2^{j}y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}(y)\right|\left|\Delta_{j}f(x-y)\chi_{B(z_{m},1)}(y)\right|\,dy$
$\leqslant 2^{Nj/p}\sum_{m=1}^{2^{k\log_{2}\Omega}\Omega}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k+j})}(y)\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x+z_{m},1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|$
$\leqslant C2^{k\log_{2}{\Omega}+Nj/p}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C2^{Nj/p}2^{-k}\left\||y|^{1+\log_{2}{\Omega}}\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C2^{Nj/p}2^{-k}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
for each $k\in\mathbb{N}$. Plugging these estimate into (2.3) yields (2.3) for all $j\geqslant 1$ ($j=0$ case also follows by the same argument).
(1) It follows from (2.3) immediately.
(2) We see $q=\infty$ case (The rest case is proved in a similar way). Clearly
$f^{\ast}(\lambda)\leqslant\left\|f\mid L^{\infty}\right\|$
holds for all $\lambda>0$ and for all $f\in L^{\infty}$. Thus, we have
$\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p/\theta,\infty}\right\|$ $=\sup_{\lambda>0}\lambda^{\theta/p}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)^{\theta}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)^{1-\theta}$
$\leqslant\left[\sup_{\lambda>0}\lambda^{1/p}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)\right]^{\theta}\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-\theta}$
$\leqslant\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{\theta}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-\theta}$
(2.3) $\leqslant C2^{jN(1-\theta)/p}\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|$
for all $z\in\mathbb{R}^{N}$. ∎
Remark 2.11.
Let $1<p<\infty$ and $1\leqslant q\leqslant\infty$. Then the following embeddings
$L^{p,q}_{\mathrm{ul}}\subset L^{p,\infty}_{\mathrm{ul}}\subset M^{p}_{Q}$
hold for all $1\leqslant Q<p$, where $M^{p}_{Q}$ denotes the local Morrey spaces (see Definition 0.1 in [17]). The first one is a standard fact (see Proposition 1.4.10 in [7]). The second one is obtained as follows. For fixed $z\in\mathbb{R}^{N}$ and $R\in\left]0,1\right]$, we find out that
$R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|$ $=R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,1)}\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|$
$\leqslant\sup_{R>0}R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,1)}\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|\left(=\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}\mathcal{M}^{p}_{Q}\right\|\right)$
[…]
Combining these embeddings with (2.4) in [17] also leads Proposition 2.10.
We use the reiteration theorem for Besov spaces in the proof of Proposition 3.2.
Proposition 2.12.
(Reiteration theorem)
Fix $1\leqslant p\leqslant\infty$, $0<\kappa<1$ and $r,r_{0},r_{1}\in\left[1,\infty\right]$. Let $s,s_{0},s_{1}$ satisfty $s=(1-\kappa)s_{0}+\kappa s_{1}$. Then we have the formula
$B^{s}_{p,r}=\left(B^{s_{0}}_{p,r_{0}},B^{s_{1}}_{p,r_{1}}\right)_{\kappa,r}.$
In addition, assume $1<p<\infty$. Then we have the formula
$B^{s}_{(p,\infty),r}=\left(B^{s_{0}}_{(p,\infty),r_{0}},B^{s_{1}}_{(p,\infty),r_{1}}\right)_{\kappa,r}.$
Proof.
See Theorem 6.4.5.(1) in [4]. For the latter, combine Theorem 2.2 in [19] and Definition 4.5 in [19], where we take $E=L^{p,\infty}_{\mathrm{ul}}$. ∎
2.4 Decay estimates of $S(t)$
In this subsection, we collect the decay estimates of operators $S(t)$ on our function spaces. We begin with Besov–Lorentz spaces.
Proposition 2.13.
Let $p\in\left]1,\infty\right[$, $r\in\left[1,\infty\right]$ and $s\geqslant\sigma$. Then, there exists $C>0$ such that the estimate
$\left\|S(t)f\mid B^{s}_{(p,\infty),r}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mid B^{\sigma}_{(p,\infty),r}\right\|$ (2.5)
holds for all $t>0$ and for all $f\in B^{\sigma}_{(p,\infty),r}$. Furthermore, if $s>\sigma$, the estimate
$\left\|S(t)f\mid B^{s}_{(p,\infty),1}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mid B^{\sigma}_{(p,\infty),\infty}\right\|$ (2.6)
holds for all $t>0$ and for all $f\in B^{\sigma}_{(p,\infty),\infty}$.
Proof.
We can prove it in the same way as Theorem 2.1 in [30] with the aid of Proposition 2.3 and 2.12 (see also [1, 18, 17]). ∎
We shall make use of the decay estimates of operators $S(t)$ on Lorentz spaces.
Proposition 2.14.
Let $1<p\leqslant q<\infty$. Then, there exists $C>0$ such that the estimate
$\left\|S(t)f\mid L^{q,\infty}_{\mathrm{ul}}\right\|\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|$. (2.7)
holds for all $t>0$ and for all $f\in L^{p,\infty}_{\mathrm{ul}}$.
Proof.
$p=q$ case. By using Proposition 2.3 and Lemma 2.1 in [21], we gain
$\left\|S(t)f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\ast f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$=\left\|\mathcal{F}^{-1}\left[\exp\left(-|\xi|^{\theta}\right)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|=C\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$.
$p<q$ case. By Propositions 2.9, 2.10 and 2.13, it is seen that
$\left\|S(t)f\mathrel{}\middle|\mathrel{}L^{q,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{0}_{(q,\infty),1}\right\|=C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{\frac{N(q-p)}{qp}-\frac{N(1-p/q)}{p}}_{\left(\frac{p}{p/q},\infty\right),1}\right\|$
$\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{\frac{N(q-p)}{qp}}_{(p,\infty),1}\right\|\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),\infty}\right\|$
$\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
holds for all $t>0$ and for all $f\in L^{p,\infty}_{\mathrm{ul}}$. ∎
Remark 2.15.
One cannot drop “1+1+” in (2.7); one arrives at a contradiction by considering $f(x)\equiv 1$ and taking the limit $t\to\infty$, for instance. This is a “cost” of employing uniformly local function spaces. In this article, however, the term “1+1+” is negligible, as we restrict our attention to local-in-time solvability and take $t<1$.
2.5 An estimate on the forcing term
In this subsection, we establish the following proposition that will be used in the proof of Theorem 1.9; therein lies the novelty of this paper.
Proposition 2.16.
Let $T>0$ and $p\in\left]1,\infty\right[$. There exists a poitive constant $C$ depending on $T$ such that the inequality
[…]
To the best of the author’s knowledge, this estimate seems not to have received attention, even for usual Besov spaces and for $\theta=2$. We note that the proof below does not depend on the base space $L^{p,\infty}_{\mathrm{ul}}$ except for the convolution inequality of Young type (Proposition 2.3).
Proof of Proposition 2.16.
We set
$C_{T}(\xi)=C_{T}^{\theta}(\xi):=\left(\int_{0}^{T}\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\,dt\right)^{-1}=\frac{|\xi|^{2\theta}}{|\xi|^{\theta}T+e^{-|\xi|^{\theta}T}-1}$
and observe that
$C_{T/2^{\theta j}}(\xi)=2^{2\theta j}C_{T}(2^{-j}\xi)$ (2.8)
holds for all $T>0$, $j\in\mathbb{Z}$ and $\xi\in\mathbb{R}^{N}\setminus\{0\}$. We calculate as follows
$\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right](\xi)$ $=\int_{0}^{t}e^{-(t-\tau)|\xi|^{\theta}}\mathcal{F}\mu(\xi)\,d\tau=\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\mathcal{F}\mu(\xi).$ (2.9)
Then we obtain
$\mathcal{F}\Delta_{j}\mu$ $=\varphi_{j}(\xi)\mathcal{F}\mu(\xi)\times 1=\varphi_{j}(\xi)\mathcal{F}\mu(\xi)C_{T}(\xi)\int_{0}^{T}\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\,dt$
(2.9) $=\Phi_{j}(\xi)C_{T}(\xi)\int_{0}^{T}\varphi_{j}(\xi)\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right](\xi)\,dt$
for all $j\in\mathbb{Z}_{\geqslant 1}$. Here $\Phi_{j}$ is the same as in the proof of Proposition 2.10. Thanks to Proposition 2.3, we see that
$\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\ast\int_{0}^{T}\Delta_{j}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right]\,dt\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T}\left\|\Delta_{j}\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,dt$
holds for all $T>0$. Now we replace $T$ with $T/2^{\theta j}$ to get
$\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T/2^{\theta j}}\left\|\Delta_{j}\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|dt$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$\qquad\qquad\qquad\times\int_{0}^{T/2^{\theta j}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|\,dt.$ (2.10)
Note that we have
$C^{\sharp}_{T}:=T\int_{\mathbb{R}^{N}}\left|\frac{1}{(2\pi)^{N}}\int_{\text{supp}\Phi_{0}}\Phi_{0}(\xi)C_{T}(\xi)e^{\sqrt{-1}x\xi}\,d\xi\right|\,dx<+\infty$
because $\xi\mapsto\Phi_{0}(\xi)C_{T}(\xi)$ is a smooth function. Thus, (2.8) gives
$\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|$ $=2^{2\theta j}\left\|\mathcal{F}^{-1}\left[\Phi_{0}(2^{-j}\xi)C_{T}(2^{-j}\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$=2^{2\theta j}\left\|\mathcal{F}^{-1}\left[\Phi_{0}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|=2^{2\theta j}T^{-1}C^{\sharp}_{T}.$ (2.11)
Since the indices $j$ are non-negative, we deduce that
$\int_{0}^{\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|dt$ $\leqslant\frac{T}{2^{\theta j}}\sup_{0\leqslant t\leqslant\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|$
$(j\geqslant 0\Rightarrow)\quad\leqslant\,\frac{T}{2^{\theta j}}$ $\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$ (2.12)
Plugging (2.5) and (2.5) into (2.5) yields
$2^{-\theta j}\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C^{\sharp}_{T}\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|$
for all $j\in\mathbb{Z}_{\geqslant 1}$. In a similar way, we can estimate $\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ as follows:
$\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant T\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$
From Appendix A, we have
$T^{-1}C^{\flat}_{T}:=\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mid L^{1}\right\|<\infty.$ (2.13)
Therefore, taking the supremum of $2^{-\theta j}\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ over $j\in\mathbb{N}$ implies
$\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|$ $\leqslant\left(C^{\flat}_{T}+C^{\sharp}_{T}\right)\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$
Finally we invoke Proposition 2.9 to reach the desired inequality. ∎
Remark 2.17.
  1. 1.
    Because $C_{T}(\xi)$ has an exceptional point ($\xi=0$), the smoothness $\Phi_{(0)}(\xi)C_{T}(\xi)$ is not trivial. In fact, we can confirm that the one-dimensional function
    $\mathbb{R}\ni x\mapsto\frac{x^{2}}{x+e^{-x}-1}=:\Psi(x)\in\mathbb{R}$
    belongs to $C^{\infty}(\left[0,\infty\right[)$. Thus it is all okay for $\theta=2$ case, for the map $\xi\mapsto|\xi|^{2}$ is smooth in $\mathbb{R}_{\xi}^{N}$ and $C_{T}(\xi)=T^{-2}\Psi(T|\xi|^{2})$ , but the rest cases require additional care (see Appendix A).
  2. 2.
    Taking into account (2.5), we cannot obtain homogeneous Besov-type space version of the inequality (see also (A)). This is consistent with Theorem 1.5 (1) (we note that $B^{-\theta}_{(p,\infty),1}\setminus\dot{B}^{-\theta}_{(p,\infty),\infty}\neq\varnothing$).
  3. 3.
    The coefficient $C^{\flat}_{T}+C^{\sharp}_{T}$ appering in the proof of Proposition 2.16 satisfies
    $0<C^{\flat}_{T}+C^{\sharp}_{T}\leqslant CT^{-1}\quad\text{as }T\to 0.$
3 Proofs of main theorems
In this section, we show the existence of solution to (1.2) as a fixed point of the map defined below in Lorentz spaces. We carry out partially the same arguments as in [11], but we give a full proof for reader’s convenience.
To begin with, we fix $T>0$ and $p\in\left]1,\infty\right[$ and we write
$X_{T}:=\left\{u(x,t):\text{Lebesgue measurable in }\mathbb{R}^{N}\times\left]0,T\right[\mathrel{}\middle|\mathrel{}\left\|u\mid X_{T}\right\|<\infty\right\}$
, where
$\left\|u\mid X_{T}\right\|:=\sup_{0<t<T}\left\|u(\cdot,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|$.
For $u\in X_{T}$ and $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$, we define operators $I$, $J$ and $\mathcal{K}$ as follows:
$\mathcal{K}[u]=\mathcal{K}[u](x,t;\mu)$ $=I[\mu](x,t)+J[u](x,t)$
$:=\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau+\int_{0}^{t}S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right](x)\,d\tau$
$=\int_{0}^{t}S(\tau)\mu(x)\,d\tau+\int_{0}^{t}S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau.$
We take a constant $M>0$ (which is determined later by (3.2)) and set
$B_{M}:=\left\{u\in X_{T}\mid\left\|u\mid X_{T}\right\|\leqslant M\right\}.$
Then $\left(B_{M},\left\|\cdot-\cdot\mathrel{}\middle|\mathrel{}X_{T}\right\|\right)$ is a complete metric space. In subsequent subsections, it turns out that the map $\mathcal{K}:B_{M}\to B_{M}$ is well-defined and is a contraction map.
3.1 Estimates on $I[\mu]$
In this subsection, we show two estimates on the operator $I[\mu]$ which correspond to Theorem 1.5 (2) and 1.5 (1). These kind of critical estimates were early obtained independently by Meyer [20] and Yamazaki [26] in Lorentz spaces. Recently, in homogeneous Besov spaces $\dot{B}^{s}_{p,\infty}$ [6] or in uniformly local weak Zygmund spaces [11], such critical estimate was obtained and played a key role.
Proposition 3.1.
Fix $0<T<1$, $1<p<\infty$ and $r\in\left]1,p\right[$. Then $I[\mu]$ is a map from $B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}$ to $X_{T}$, and its operator norm is finite.
Proof.
We define $k(r)>0$ as $p^{-1}=(2r)^{-1}+(2k(r))^{-1}$. It is easy to see that $1<r<p<k(r)$. Then we let $\lambda>0$, set
$\eta=\eta(\lambda):=\lambda^{a},\quad a:=\frac{\theta pr}{2N(p-r)}$ (3.1)
and decompose $I[\mu]$ as
$I[\mu](x,t)$ $=\int_{0}^{t}S(\tau)\mu(x)\,d\tau=\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\mu(x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\mu(x)\,d\tau=:I_{\eta}^{0}[\mu](x,t)+I_{\eta}^{1}[\mu](x,t).$
By Proposition 2.9, Proposition 2.13 and (3.1), we have
$\left\|I_{\eta}^{0}[\mu](x,t)\mid L^{r,\infty}_{\mathrm{ul}}\right\|\leqslant C\int_{0}^{\eta}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0}_{(r,\infty),1}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\tau^{-1+\frac{N}{\theta r}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|=C\lambda^{1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|$
and
$\left\|I_{\eta}^{1}[\mu](x,t)\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\leqslant C\int_{\eta}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\mu(x)\mathrel{}\middle|\mathrel{}B^{0}_{(k(r),\infty),1}\right\|\,d\tau$
$\leqslant C\int_{\eta}^{\infty}\tau^{-1+\frac{N}{\theta k(r)}-\frac{N}{\theta p}}\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|=C\lambda^{-1/2}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|.$
Now we use Corollary 2.8 in the form $L^{p,\infty}_{\mathrm{ul}}\supset\left(L^{r,\infty}_{\mathrm{ul}},L^{k(r),\infty}_{\mathrm{ul}}\right)_{1/2,\infty}$ to derive
$\left\|I[\mu](x,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{I[\mu]=I_{0}+I_{1}}\left\{\left\|I_{0}\mathrel{}\middle|\mathrel{}L^{r,\infty}_{\mathrm{ul}}\right\|+\lambda\left\|I_{1}\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|I_{\eta(\lambda)}^{0}[\mu]\mathrel{}\middle|\mathrel{}L^{r,\infty}_{\mathrm{ul}}\right\|+\lambda^{1/2}\left\|I_{\eta(\lambda)}^{1}[\mu]\mathrel{}\middle|\mathrel{}L^{k(r),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\left\{\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|+\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{k(r)}-\frac{N}{p}}_{(k(r),\infty),\infty}\right\|\right\}$
$\leqslant C\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta+\frac{N}{r}-\frac{N}{p}}_{(r,\infty),\infty}\right\|$
for all $t<T$. In the last inequality, we used Proposition 2.10. ∎
Proposition 3.2.
Fix $0<T<1$. Let $1<p<\infty$. Then $I[\mu]$ is a map from $B^{-\theta}_{(p,\infty),1}$ to $X_{T}$, and its operator norm is finite.
Proof.
We set
$\Psi_{j}(x):=\mathcal{F}^{-1}\left[\Phi_{j}\right](x)\,\,\text{for }j\in\mathbb{Z},\quad\Psi_{(0)}(x):=\mathcal{F}^{-1}\left[\Phi_{(0)}\right](x)$
, where $\Phi_{j}$ and $\Phi_{(0)}$ are the same as in the proof of Proposition 2.16. By using Proposition 2.3 and 2.9, we have
$\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant C\left\|\int_{0}^{t}S(\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),1}\right\|$
$=C\sum_{j\in\mathbb{N}}2^{0j}\left\|\Delta_{j}\int_{0}^{t}S(\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$=C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\ast\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C\sum_{j\in\mathbb{N}}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$. (3.2)
Fix $j\in\mathbb{N}$ and sufficiently small $r>0$. Because $\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}L^{1}\right\|$ is independent of $j\in\mathbb{N}$, Proposition 2.12 and 2.13 give that
$\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|\leqslant\sum_{l=-2}^{2}2^{\theta(j+l)}2^{-\theta(j+l)}\left\|\Delta_{j+l}\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$\leqslant 5\cdot 2^{2\theta}2^{-\theta j}\max_{-2\leqslant l\leqslant 2}2^{\theta(j+l)}\left\|\Delta_{j+l}\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$\leqslant C2^{-\theta j}\left\|\int_{0}^{t}S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}B^{\theta}_{1,\infty}\right\|$
$=C2^{-\theta j}\left\|\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[\Psi_{j}\right](x)\,d\tau\mathrel{}\middle|\mathrel{}\left(B^{\theta-\theta r}_{1,\infty},B^{\theta+\theta r}_{1,\infty}\right)_{1/2,\infty}\right\|$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\chi_{\left\{0<\tau<t\right\}}(\tau)\left(1+\tau^{r-1}\right)\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}B^{0}_{1,\infty}\right\|\,d\tau$
$\qquad\qquad\qquad\qquad\quad+\lambda\int_{\lambda^{1/2r}}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left(1+\tau^{-1-r}\right)\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}B^{0}_{1,\infty}\right\|\,d\tau\bigg\}$
$\leqslant C2^{-\theta j}\sup_{\lambda>0}\lambda^{-1/2}\bigg\{\int_{0}^{\lambda^{1/2r}}\tau^{r-1}\,d\tau+\lambda\int_{\lambda^{1/2r}}^{\infty}\tau^{-1-r}\,d\tau\bigg\}\left\|\Psi_{j}(x)\mathrel{}\middle|\mathrel{}L^{1}\right\|=C2^{-\theta j}$
holds for $0<t\leqslant T<1$ (Here we wrote $\Delta_{j}:=\Delta_{0}$ for $j<0$). Plugging this estimate in (3.1) and taking the supremum over $t\in\left]0,T\right[$ yield the claim. ∎
3.2 Contraction mapping argument
In this section, we deal with $J[u]$ and complete the proof of main theorems.
Lemma 3.3.
Assume that $\mu$ satisfies $I[\mu]\in B_{M/2}$ and $p$ satisfies $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Then, the operator $u\mapsto\mathcal{K}[u]$ is a map from $B_{M}$ to $B_{M}$.
In the following proof, we divide the case into $p>\frac{N(\gamma-1)}{\theta}$ and $p=\frac{N(\gamma-1)}{\theta}$. While they can be discussed jointly with minor adjustments, we separate them here for simplicity.
Proof.
$p>\frac{N(\gamma-1)}{\theta}$ case. Take arbitrary $u\in B_{M}$. By utilizing Proposition 2.14 and Lemma 2.1 (2), we have
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\right\}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant Ct^{1-\frac{N(\gamma-1)}{\theta p}}\left\|u\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant CT^{1-\frac{N(\gamma-1)}{\theta p}}M^{\gamma}\leqslant M/2$
for all $0<t<T\leqslant 1$. This gives $\mathcal{K}[u]\in B_{M}$.
$p=\frac{N(\gamma-1)}{\theta}$ case. Pick arbitrary $u\in B_{M}$. We take $\rho>0$ and $k(\rho)>0$ such that
$1<p/\gamma<\rho<p=\frac{N(\gamma-1)}{\theta}<k(\rho)\quad\text{and}\quad\frac{1}{p}=\frac{1/2}{\rho}+\frac{1/2}{k(\rho)}.$ (3.3)
Then we let $\lambda>0$, set
$\eta=\eta(\lambda):=\lambda^{b},\quad b:=\frac{\theta p\rho}{2(\theta p\rho+Np-N\gamma\rho)}$ (3.4)
and decompose $J[u]$ as
$J[u](x,t)$ $=\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=:J_{\eta}^{0}[u](x,t)+J_{\eta}^{1}[u](x,t).$
By Proposition 2.14, (3.3) and (3.4), we have
$\left\|J_{\eta}^{0}[u](x,t)\mid L^{\rho,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\int_{0}^{\eta}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\mid L^{\rho,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\left(1+\tau^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)\left\||u|^{\gamma-1}u(\cdot,t-\tau)\mid L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)+1}\left\|u\mid L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant C\eta^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)+1}M^{\gamma}$
$=C\lambda^{1/2}M^{\gamma}$
and
$\left\|J_{\eta}^{1}[u](x,t)\mid L^{k(\rho),\infty}_{\mathrm{ul}}\right\|$ $\leqslant\int_{\eta}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\mid L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{\eta}^{\infty}\tau^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)}\left\||u|^{\gamma-1}u(\cdot,t-\tau)\mid L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)+1}\left\|u\mid L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant C\eta^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)+1}M^{\gamma}$
$=C\lambda^{-1/2}M^{\gamma}.$
The use of Corollary 2.8 in the form $L^{p,\infty}_{\mathrm{ul}}\supset\left(L^{\rho,\infty}_{\mathrm{ul}},L^{k(\rho),\infty}_{\mathrm{ul}}\right)_{1/2,\infty}$ implies that
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{J[\mu]=J_{0}+J_{1}}\left\{\left\|J_{0}\mathrel{}\middle|\mathrel{}L^{\rho,\infty}_{\mathrm{ul}}\right\|+\lambda\left\|J_{1}\mathrel{}\middle|\mathrel{}L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|J_{\eta(\lambda)}^{0}[u]\mathrel{}\middle|\mathrel{}L^{\rho,\infty}_{\mathrm{ul}}\right\|+\lambda^{1/2}\left\|J_{\eta(\lambda)}^{1}[u]\mathrel{}\middle|\mathrel{}L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant CM^{\gamma}\leqslant M/2.$ (3.5)
Thus we have
$\sup_{0<t<T}\left\|J[u](\cdot,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\frac{M}{2}$
and therefore we find out $\mathcal{K}[u]\in B_{M}$. ∎
Lemma 3.4.
Assume that $p$ satisfies the same conditions as in Lemma 3.3. Then, the map $\mathcal{K}:B_{M}\to B_{M}$ is a contraction map.
Proof.
Firstly by using Lemma 2.1, we prepare the following estimate. For all $q>1$, $\sigma\geqslant 1$ and for all $f$, $g\in L^{q\sigma,\infty}_{\mathrm{ul}},$ we have
$\left\||f|^{q-1}f-|g|^{q-1}g\mathrel{}\middle|\mathrel{}L^{\sigma,\infty}_{\mathrm{ul}}\right\|$ $\leqslant q\left\|\max\{|f|;|g|\}^{q-1}|f-g|\mathrel{}\middle|\mathrel{}L^{\sigma,\infty}_{\mathrm{ul}}\right\|$
$\leqslant q\left\|\max\{|f|;|g|\}^{q-1}\mathrel{}\middle|\mathrel{}L^{\frac{q\sigma}{q-1},\infty}_{\mathrm{ul}}\right\|\left\|f-g\mid L^{q\sigma,\infty}_{\mathrm{ul}}\right\|$
$\leqslant q\left(\left\|f\mathrel{}\middle|\mathrel{}L^{q\sigma,\infty}_{\mathrm{ul}}\right\|^{q-1}+\left\|g\mathrel{}\middle|\mathrel{}L^{q\sigma,\infty}_{\mathrm{ul}}\right\|^{q-1}\right)\left\|f-g\mid L^{q\sigma,\infty}_{\mathrm{ul}}\right\|$. (3.6)
$p>\frac{N(\gamma-1)}{\theta}$ case. Take any $u,v\in B_{M}$. For all $0<t<T\leqslant 1$, it holds that
$\left\|\mathcal{K}[u](\cdot,t;\mu)-\mathcal{K}[v](\cdot,t;\mu)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|=\left\|J[u](\cdot,t)-J[v](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\quad\leqslant C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)-|v|^{\gamma-1}v(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\right\}\left\||u|^{\gamma-1}u(\cdot,\tau)-|v|^{\gamma-1}v(\cdot,\tau)\mathrel{}\middle|\mathrel{}L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant C\int_{0}^{t}(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\left(\left\|u\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma-1}+\left\|v\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma-1}\right)\left\|u-v\mid L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant Ct^{1-\frac{N(\gamma-1)}{\theta p}}M^{\gamma-1}\left\|u-v\mid X_{T}\right\|\leqslant\frac{1}{2}\left\|u-v\mid X_{T}\right\|$
through the use of (3.2) and Proposition 2.14.
$p=\frac{N(\gamma-1)}{\theta}$ case. Utilizing (3.2) and carrying out a similar procedure as in the proof of Lemma 3.3, the claim follows. ∎
Proof of Theorem 1.5.
From Lemma 3.3, Lemma 3.4 and the contraction mapping theorem, we only have to show $\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|\leqslant M/2$. Assertion (1) is proved by Proposition 3.2. We next prove (3). This is established by following calculation. We use Proposition 2.9 and 2.13 to see that
$\left\|I[\mu](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),1}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{s/\theta}\right\}\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\,d\tau$
$\leqslant C\left(t+t^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|$
holds for all $0<t\leqslant T$ and thus
$\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|$ $\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\leqslant\frac{M}{2}$ (3.7)
holds for sufficiently small $T>0$. Finally we prove (2). Assume that $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfies
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|<\delta.$
$\delta>0$ is determined later. By definition of $\limsup$, there exists an integer $m\in\mathbb{Z}$ such that $2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|<\delta$ holds for all $j\geqslant m$. Set $\mu_{1}:=\sum_{j=0}^{m}\Delta_{j}\mu$ and $\mu_{2}:=\mu-\mu_{1}$. Then for every $j\in\mathbb{N}$, we have the following relations.
$\Delta_{j}\mu_{1}$ $=\begin{cases}\Delta_{j}\mu&\text{for }j\leqslant m-1\\ \left(\Delta_{m-1}+\Delta_{m}\right)\Delta_{j}\mu&\text{for }j=m,m+1\\ 0&\text{for }j\geqslant m+2,\end{cases}$
$\Delta_{j}\mu_{2}$ $=\begin{cases}0&\text{for }j\leqslant m-1\\ \left(\Delta_{m+1}+\Delta_{m+2}\right)\Delta_{j}\mu&\text{for }j=m,m+1\\ \Delta_{j}\mu&\text{for }j\geqslant m+2.\end{cases}$
Thus we have $\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|<\delta$ and
$\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma}_{(p,\infty),\infty}\right\|$ $=\max_{0\leqslant j\leqslant m+1}2^{\sigma j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\max_{0\leqslant j\leqslant m+1}2^{(\sigma+\theta)j}2^{-\theta j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant 2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|\leqslant 2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|$
for some $\sigma\in\left]-\theta,0\right[$. Now applying (3.7) and Proposition 3.1 to $\mu_{1}$ and $\mu_{2}$ respectively, we get
$\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|$ $\leqslant\left\|I[\mu_{1}]\mathrel{}\middle|\mathrel{}X_{T}\right\|+\left\|I[\mu_{2}]\mathrel{}\middle|\mathrel{}X_{T}\right\|$
$\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma}_{(p,\infty),\infty}\right\|+C_{\varepsilon}\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{(\frac{Np}{N+p\varepsilon},\infty),\infty}\right\|$
$\leqslant C\left(T+T^{1+s/\theta}\right)2^{\theta m}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{(\frac{Np}{N+p\varepsilon},\infty),\infty}\right\|+C_{\varepsilon}\delta.$
Taking $\delta>0$ and $T>0$ sufficiently small, we obtain $\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|\leqslant M/2$. ∎
Remark 3.5.
Let $p$ satisfy the same conditions in Theorem 1.5. If the forcing term $\mu$ satisfies one of the conditions in Theorem 1.5, we find a number $s<0$ such that $\mu$ belongs to $B^{s}_{(p,\infty),\infty}\subset B^{s-N/p}_{\infty,\infty}$. Then the solution $u(x,t)$ constructed above satisfy the initial condition in the following sense: $u(x,t)\to 0$ in the weak-$\ast$ topology of the space $B^{s-N/p}_{\infty,\infty}$ as $t\to+0$. Take arbitrary $\psi\in B^{N/p-s}_{1,1}$ and calculate as follows:
$\left|\langle u(\cdot,t),\psi\rangle\right|$ $\leqslant\int_{0}^{t}\left|\langle S(t-\tau)\mu,\psi\rangle\right|\,d\tau+\int_{0}^{t}\left|\langle S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right],\psi\rangle\right|\,d\tau$
$\leqslant\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s-N/p}_{\infty,\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\,d\tau$
$\quad+\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{s-N/p}_{\infty,\infty}\right\|\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\,d\tau$
$\leqslant\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\,d\tau\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|$
$\quad+\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}B^{\max\left\{0;s+\frac{(\gamma-1)N}{p}\right\}}_{(p/\gamma,\infty),\infty}\right\|\,d\tau\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\bigg(\int_{0}^{t}1\,d\tau\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|$
$\qquad+\int_{0}^{t}(t-\tau)^{\min\left\{0;-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}B^{0}_{(p/\gamma,\infty),\infty}\right\|\,d\tau\bigg)$
$\leqslant C\left\|\psi\mathrel{}\middle|\mathrel{}B^{N/p-s}_{1,1}\right\|\big(\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|+\left\|u\mathrel{}\middle|\mathrel{}X_{T}\right\|^{\gamma}\big)t^{\min\left\{1;1-\frac{s}{\theta}-\frac{(\gamma-1)N}{p\theta}\right\}}.$
Conditions $p\geqslant\frac{N(\gamma-1)}{\theta}$ and $s<0$ imply that the last term disappears in the limit $t\to 0$.
Proof of Theorem 1.9.
Let $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ satisfies the integral equation (1.2) for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$, i.e., we have
$u(x,t)=I[\mu](x,t)+J[u](x,t)$
for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. The same calculation of the proof of Lemma 3.3 (estimate on $J[u]$) claims that $J[u](x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ and thus we deduce that
$I[\mu](x,t)=u(x,t)-J[u](x,t)$ $\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})).$
Finally we apply Proposition 2.16 to derive $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|<\infty$. ∎
Acknowledgements.
The author would like to thank Professor Ryo Takada for several comments. The author was supported in part by JST SPRING, Grant Number JPMJSP2108, and KAKENHI #24KJ0623.
Appendix A (2.13) for general $0<\theta\leqslant 2$
In this section, we confirm that
$\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mid L^{1}\right\|<\infty$ (A.1)
is valid for all $\theta\in\left]0,2\right]$, where
$C_{T}(\xi)=C_{T}^{\theta}(\xi):=\left(\int_{0}^{T}\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\,dt\right)^{-1}=\frac{|\xi|^{2\theta}}{|\xi|^{\theta}T+e^{-|\xi|^{\theta}T}-1}.$
We modify the argument of Lemma 2.1 in [21].
Proof.
We first note that there exists $C>0$ such that
$\left|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right](x)\right|\leqslant C\quad\text{for all }x\in\mathbb{R}^{N}$ (A.2)
because $\Phi_{(0)}(\xi)C_{T}(\xi)$ is integrable with respect to $\xi$. Thus we focus on $|x|\gg 1$.
We define the operator
$L(x,D):=\frac{xD_{\xi}}{\sqrt{-1}|x|^{2}}.$
Then we have
$L(x,D)e^{\sqrt{-1}x\xi}=\frac{1}{\sqrt{-1}|x|^{2}}\sum_{k=1}^{N}x_{k}\left(\frac{\partial}{\partial\xi_{k}}e^{\sqrt{-1}x\xi}\right)=e^{\sqrt{-1}x\xi}.$
The conjugate operator is
$L^{\ast}(x,D):=\frac{\sqrt{-1}xD_{\xi}}{|x|^{2}}.$
Thus we can write $\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]$ as
$\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}L^{\ast}(x,D)\left(\frac{|\xi|^{2\theta}\Phi_{(0)}(\xi)}{|\xi|^{\theta}T+e^{-|\xi|^{\theta}T}-1}\right)\,d\xi.$
Set a $C_{c}^{\infty}(\mathbb{R}^{N})$-function $\rho$ satisfying
$\rho(\xi)=\begin{cases}1,\quad|\xi|\leqslant 1,\\ 0,\quad|\xi|>2\end{cases}$
and decompose $\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]$ as
$\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]$
$=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}\rho\left(\frac{\xi}{\delta}\right)L^{\ast}(x,D)\left(\Phi_{(0)}(\xi)C_{T}(\xi)\right)\,d\xi$
$\quad+\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}\left[1-\rho\left(\frac{\xi}{\delta}\right)\right]L^{\ast}(x,D)\left(\Phi_{(0)}(\xi)C_{T}(\xi)\right)\,d\xi$
$=:I+II,$
where $\delta\in\left]0,1/2\right[$ to be fixed later. Recall that the function $\Psi(s):=\frac{s^{2}}{s+e^{-s}-1}$ belongs to $C^{\infty}(\left[0,\infty\right[)$. By mathematical induction on $k$, we can show that there exist constants $C^{k}_{j,i}$ which are independent of $T$ such that
$\left(L^{\ast}\right)^{k}C_{T}(\xi)$ $=\left(L^{\ast}\right)^{k}\left(\frac{|\xi|^{2\theta}}{|\xi|^{\theta}T+e^{-|\xi|^{\theta}T}-1}\right)=T^{-2}\left(L^{\ast}\right)^{k}\left(\Psi(|\xi|^{\theta}T)\right)$
$=\frac{\sqrt{-1}^{k}}{|x|^{2k}}\sum_{j=1}^{k}T^{j-2}\Psi^{(j)}(|\xi|^{\theta}T)\sum_{\begin{subarray}{c}0\leqslant i\leqslant k-j\\ 2i\leqslant k\end{subarray}}C^{k}_{j,i}|\xi|^{j\theta+2i-2k}(x\xi)^{k-2i}|x|^{2i}$
holds for all $k\geqslant 1$ and for all $\xi$. Note that we know $\left|\frac{d^{j}}{ds^{j}}\Psi\left(Ts\right)\right|\leqslant T^{j}C$ for bounded $s$ and for $j\in\mathbb{N}$. Thus we have the estimate
$\left|\left(L^{\ast}\right)^{k}C_{T}(\xi)\right|$ $\leqslant\frac{C_{k}}{|x|^{k}}\sum_{j=1}^{k}T^{j-2}|\xi|^{j\theta-k}\leqslant C_{k}T^{-1}|x|^{-k}|\xi|^{\theta-k}.$
For $I$, we have
$|I|$ $\leqslant C\left|\int_{\mathbb{R}^{N}}e^{\sqrt{-1}x\xi}\rho\left(\frac{\xi}{\delta}\right)L^{\ast}(x,D)\left(C_{T}(\xi)\Phi_{(0)}(\xi)\right)\,d\xi\right|$
$=C\int_{B_{2\delta}(0)}\left|e^{\sqrt{-1}x\xi}\right|\left|L^{\ast}(x,D)C_{T}(\xi)\right|\,d\xi$
$\leqslant C\int_{B_{2\delta}(0)}C_{1}T^{-1}|x|^{-1}|\xi|^{\theta-1}\,d\xi=CT^{-1}|x|^{-1}\delta^{\theta-1+N}.$
Next we estimate $II$. Note that the estimate
$\left|L^{\ast}(x,D)^{k}\left[1-\rho\left(\frac{\xi}{\delta}\right)\right]\right|\leqslant\frac{C_{N,k}}{\delta^{k}|x|^{k}}$
holds for $k\geqslant 1$. To estimate $II$, we integrate by parts $N$ times.
$|II|$ $\leqslant C\int_{\mathbb{R}^{N}}\left|e^{\sqrt{-1}x\xi}\right|\left|\left(L^{\ast}\right)^{N}\left\{\left[1-\rho\left(\frac{\xi}{\delta}\right)\right]L^{\ast}\left(C_{T}(\xi)\Phi_{(0)}(\xi)\right)\right\}\right|\,d\xi$
$\leqslant C\int_{\mathbb{R}^{N}\setminus B_{\delta}(0)}\left[1-\rho\left(\frac{\xi}{\delta}\right)\right]\left|\left(L^{\ast}\right)^{N+1}\left(C_{T}(\xi)\Phi_{(0)}(\xi)\right)\right|\,d\xi$
$\quad+C\int_{B_{2\delta}(0)\setminus B_{\delta}(0)}\sum_{k=1}^{N}\left|\left(L^{\ast}\right)^{k}\left[1-\rho\left(\frac{\xi}{\delta}\right)\right]\left(L^{\ast}\right)^{N-k+1}\left(C_{T}(\xi)\Phi_{(0)}(\xi)\right)\right|\,d\xi$
$=:II_{1}+II_{2}.$
Because $B_{3}(0)\subset\left\{\xi\in\mathbb{R}^{N}\mid\Phi_{(0)}(\xi)\equiv 1\right\}$ holds, it is seen that
$|II_{1}|$ $\leqslant C\left(\int_{\text{supp}\Phi_{(0)}\setminus B_{3}(0)}+\int_{B_{3}(0)\setminus B_{\delta}(0)}\right)\left|\left(L^{\ast}\right)^{N+1}\left(C_{T}(\xi)\Phi_{(0)}(\xi)\right)\right|\,d\xi$
$\leqslant\frac{C_{N+1}}{|x|^{N+1}}\int_{\text{supp}\Phi_{(0)}\setminus B_{3}(0)}T^{-2}\max_{\xi}\left|D_{\xi}^{N+1}\left[\Psi(T|\xi|^{\theta})\Phi_{(0)}(\xi)\right]\right|\,d\xi$ (A.3)
$\quad+\int_{B_{3}(0)\setminus B_{\delta}(0)}\left|\left(L^{\ast}\right)^{N+1}C_{T}(\xi)\right|\,d\xi$
$\leqslant\frac{C}{T^{2}|x|^{N+1}}+\frac{C}{T|x|^{N+1}}\int_{B_{3}(0)\setminus B_{\delta}(0)}|\xi|^{\theta-N-1}\,d\xi\leqslant\frac{C}{T^{2}|x|^{N+1}}+\frac{C\delta^{\theta-1}}{T|x|^{N+1}}$
for $II_{1}$. For $II_{2}$, we have
$|II_{2}|$ $\leqslant\sum_{k=1}^{N}\frac{C}{T\delta^{k}|x|^{k}}\int_{B_{2\delta}(0)\setminus B_{\delta}(0)}|x|^{-(N-k+1)}|\xi|^{\theta-(N-k+1)}\,d\xi=CT^{-1}|x|^{-N-1}\delta^{\theta-1}.$
Now we choose $\delta=|x|^{-1}$. Considering (A.2), we obtain the following estimate
$\left|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right](x)\right|\leqslant\frac{CT^{-1}}{\left(1+|x|\right)^{N+\theta}}+\frac{CT^{-2}}{\left(1+|x|\right)^{N+1}}\leqslant\frac{CT^{-2}}{\left(1+|x|\right)^{N+\min\{1;\theta\}}}.$
Thus we conclude that (2.13) holds for general $0<\theta\leqslant 2$. ∎
Remark A.1 (On (A)).
Here, we used the estimate
$\left|D_{\xi}^{N+1}\left[\Psi(T|\xi|^{\theta})\Phi_{(0)}(\xi)\right]\right|\leqslant\max_{\xi\in\mathrm{supp}\Phi_{(0)}\setminus B_{3}(0)}\left|D_{\xi}^{N+1}\left[\Psi(T|\xi|^{\theta})\Phi_{(0)}(\xi)\right]\right|\leqslant C_{N}<\infty.$
This attempt fails if we try to do so with $\Psi(T|\xi|^{\theta})\Phi_{j}(\xi)$ for all $j\in\mathbb{Z}$.